OpenStudy (anonymous):

If sinA=5/6, where angle A is in Quadrant 1, find the exact value of sin (1/2A)

OpenStudy (anonymous):

you know the formula for \(\sin(\frac{A}{2})\) ?

OpenStudy (anonymous):

if i recall it is \[\sin(\frac{A}{2})=\pm\sqrt{\frac{1-\cos(A)}{2}}\]

OpenStudy (anonymous):

so what you need is \(\cos(A)\) and you can then plug that number in

OpenStudy (anonymous):

yes, ok i got that part

OpenStudy (anonymous):

i tried doing the problem myself, may you please tell me if this answer is correct? Sin1/2A = radical6-radical11/12